OpenStudy (anonymous):

how do you do elimination using multiplication ? example problem : 6x+y=-39 3x+2y=-15

hero (hero):

6x + y = -39 3x + 2y = -15 Multiply the second equation by -2 to get: 6x + y = -39 -6x - 4y = 30 -3y = -9 Continue solving for y

OpenStudy (anonymous):

Thanks .